state and prove the necessary conditions for the existence of maximum value and minimum value?
1
Expert's answer
2021-09-16T07:49:54-0400

It is said that a function has at a point c maximum or minimum if there exists such number "\delta>0" that for all "x\ \epsilon(c-\delta,c+\delta)" we have that f(x)"\le f(c)" or f(x)"\ge\ f(c)" .

Sufficient Conditions for One Independent Variable

To develop criteria establishing whether a stationary point is a local maximum or minimum, we begin by performing a Taylor series expansion about the stationary point xo.

y(x) = y(xo) + y'(xo) (x - xo) + ½ y''(xo) (x - xo)2+ higher order terms

Now, select x sufficiently close to xo so the higher order terms become negligible compared to the second-order terms. Since the first derivative is zero at the stationary point, the above equation becomes y(x) = y(xo) + ½y" (xo) (x - xo)

Now we determine if xo is a local maximum or minimum by examining the value of y"(xo), since (x - xo)2 is always positive. If y"(xo) is positive, then the terms ½y"(xo) (x - xo)2 will always add to y(xo) in the above equation for x taking on values that are less than or greater than xo. For this case y(xo) is a local minimum.

This is summarized in the following:

If y''(xo) > 0 then y(xo) is a minimum

y''(xo) < 0 y(xo) is a maximum

y''(xo) = 0 no statement can be made

If the second derivative is zero, it is necessary to examine higher order derivatives. In general if y''(xo) = ... = yn-1(xo) = 0, the Taylor series expansion becomes

y(x) = y(xo) +"(\frac{1}{n!})" y(n)(xo) (x - xo)n

If n is even, then (x - xo)n is always positive, and the result is:

If y(n)(xo) > 0 then y(xo) is a minimum

y(n)(xo) < 0 y(xo) is a maximum